### v2

# Holey sheets: Double-Threshold Rupture of Draining Liquid Films

Abstract
Classical rupture is attributed to molecular (van der Waals) forces acting at nanometric thicknesses. Nonetheless, micron-thick liquid sheets routinely perforate far above the scale where these molecular forces act, yet the mechanism that selects opening versus healing has remained unclear. Using direct numerical simulations of a draining sheet with an entrained air bubble (cavity), we show that irreversible rupture occurs only when a deterministic double-threshold is crossed: (i) the outward driving (from airflow or inertia) is strong enough and (ii) the cavity is distorted enough. If either condition falls short, surface tension heals the cavity and the sheet reseals. The time for this process is set by the balance between inertia and viscosity – fast for inertia-dominated sheets and slower for viscous ones. This double-threshold mechanism explains why micrometer-thick films perforate and offers practical control options – driving strength and defect geometry – for predicting and controlling breakup in spray formation processes, wave breaking, and respiratory films.
preprint: APS/123-QED
The rupture of thinning liquid sheets is a ubiquitous route to fluid fragmentation, underlying phenomena from respiratory aerosolization to agricultural sprays and wave breaking [1]. During the recent COVID-19 pandemic [2], the public became acutely aware of how violent expiratory events such as coughing and sneezing produce virus-laden droplets. High-speed visualization of real-life coughs [3, 4, 5], and experiments with mechanical cough devices [6, 7] show that a mucosalivary liquid layer in the respiratory tract can be sheared into a thin bag-like sheet that drains and thins rapidly. Eventually, holes appear spontaneously in these thinning sheets, which expand and cause the sheet to break apart, resulting in a cloud of droplets. Similar sheet rupture and drop-formation processes occur in a variety of contexts, including ocean wave breaking [1], agricultural spray dispersal [8], rain-induced aerosols [9], spray cleaning [10], and drop impact on liquid pools [11, 12] or even solids [13] (fig. 1a). Quantifying how such droplets are generated is crucial for accurate risk assessments of airborne disease transmission [4], optimizing pesticide applications [8], and quantifying sea spray aerosol generation [14].
A key step in all of these processes is the nucleation of holes in the thinning sheet, which sets the eventual droplet size distribution [15, 16]. However, despite its ubiquity, the physical origin of these holes remains debated [17] – especially for micrometer-thick sheets, where molecular mechanisms (van der Waals forces or thermal fluctuations) are too weak to explain rupture [18]. Consistent with this view, experiments show that breakup typically occurs once sheets thin to micron scales and that holes nucleate at internal defects – entrained bubbles or oil droplets – within the sheet [19, 6, 20, 21, 11, 22]. In particular, bubbles in so-called Savart water sheets produced by a jet impinging on a disk have been observed to perforate the sheet [23, 24]. Yet the pathway by which a bubble or other impurity triggers hole formation—particularly under realistic sheet-drainage conditions—remains unresolved because the nucleation events are rapid and difficult to control and visualize.
In this Letter, we show that an air bubble entrained in a draining liquid sheet can nucleate a hole long before molecular-scale forces become relevant. Axisymmetric direct numerical simulations performed with the Navier-Stokes and combined level-set-volume-of-fluid solver Basilisk C [25, 26] isolate this impurity-driven pathway under controlled conditions while retaining realistic rim dynamics. To focus on the essential physics, we replace the inflating bag geometry of air-blast breakup by an initially flat sheet of thickness $h_{0}$ that drains radially with $u(r)=\omega r$ (fig. 1b), a template that captures both external-airflow forcing in bag breakup [38] and inertia-driven drainage after drop impact [39]. The dynamics are governed by the Ohnesorge number $Oh=\eta/\sqrt{\rho\gamma R_{0}}$, the Bond number $Bo=\rho\omega^{2}R_{0}^{3}/\gamma$, and the non-dimensionlized offset $\chi/R_{0}$ of the bubble’s center from the centerlines of the flow. Here $\eta$ and $\rho$ are the liquid viscosity and density, $\gamma$ is the surface tension, and $R_{0}$ the bubble radius. The driving acceleration $\omega^{2}R_{0}$ acts as an effective radial gravity. In addition, the initial cavity distortion is characterized by an open angle $\theta$ (fig. 1c), see supplementary material [26] for details on drainage mechanism leading to this geometry.
Our two key results are as follows: First, the entrained bubble grows a through-cavity that overcomes surface tension and nucleates a hole at micrometer-scale thicknesses, explaining rupture without invoking nanoscale physics. Second, the post-nucleation fate is deterministic and requires crossing a double-threshold for irreversible opening: the outward drainage must exceed a critical $Bo_{c}(Oh)$ and the initial cavity distortion – characterized by an opening angle $\theta$ – must exceed a critical $\theta_{c}(Oh)$. Otherwise, rims collide and the sheet heals. The opening-healing boundary collapses onto simple asymptotic scalings: $Bo_{c}=\mathcal{O}(1)$ for $Oh\ll 1$ and $Bo_{c}\sim Oh^{-2}$ for $Oh\gg 1$. These results rationalize the observed perforation of micron-thick sheets by internal defects and provide a predictive framework for impurity-induced breakup in natural and engineered fragmentation flows.
Phenomenology – We first examine sheet-centered bubbles ($\chi/R_{0}=0$) across the $(Oh,Bo)$ space. As the sheet thins, its top and bottom interfaces meet at the bubble poles and a through-cavity opens. Capillary waves invert the cavity from convex to concave, after which two outcomes are observed (fig. 2a,b): either the rims retract and the hole expands, breaking the sheet, or capillarity drives rim collision and the sheet heals. A systematic sweep in $(Oh,Bo)$ reveals a sharp transition curve (fig. 2c): strong radial driving (large $Bo$) yields opening, whereas weaker driving permits healing unless $Bo$ exceeds a viscosity-dependent threshold. The location and slope of this boundary reflect the competition between capillary-driven rim closure and drainage-driven advection, anticipating the timescale analysis that follows. We additionally show that delayed nucleation–representing chemical/thermal heterogeneity–can be encoded by a larger initial cavity characterized by an opening angle $\theta$, which shifts the transition and thereby introduces geometry as a second way of control.
Figure 1: (a) Liquid sheets characterized by control parameters OhOh and $Bo$ are relevant in several phenomena across the entire parameter space. (b) Schematic side view of a liquid sheet that is radially draining and, subsequently, thinning along the axial direction. The flow directions of the liquid are indicated in red. A bubble with radius $R_{0}$ is placed axisymmetrically but offset axially by a distance $\chi$. (c) When additional physical factors delay the nucleation of the hole, the initial conditions used in simulations are characterized by the polar angle $\theta$ made at the cavity edge.
Figure 2: As the liquid sheet drains radially and thins axially, the interfaces of the bubble and sheet merge to create a cavity. The time instants are shown for $Oh=0.1$, (a) and at small $Bo=10^{-3}$ where the sheet heals, while (b) at larger $Bo=10^{-2}$, the sheet opens up. The left panel depicts velocity magnitude $\|\bm{v}\|/V_{0}$, where $V_{0}=\sqrt{Bo\gamma/\rho R_{0}}$, and the black arrows depict the velocity direction. The right panel illustrates viscous dissipation $\xi=2Bo\left(\bm{\mathcal{D}:\mathcal{D}}\right)$, where $\bm{\mathcal{D}}=\left(\bm{\nabla u+\nabla u}^{T}\right)/2$ is the symmetric part of velocity gradient tensor. (c) The regime map in the log-log parameter space of OhBoOh-Bo. The transition lines at small OhOh are shown by a constant BoBo line in gray, while, at large OhOh, the transition is indicated by a black line with scaling $Oh\sim Bo^{-1/2}$.
Relevant timescales set the healing to opening transition – After the bubble’s and sheet’s interfaces merge and create an initial cavity through the sheet, we measure the time it takes for the two opposing rims to collide at the sheet’s central axis; we term this the “collision time” $t_{c}$. Physically, $t_{c}$ represents the healing time when the sheet successfully reseals. Fig. 3(a) shows $t_{c}$ (normalized by the inertio-capillary timescale $\tau_{\gamma}=\sqrt{\rho R_{0}^{3}/\gamma}$) as a function of $Oh$ for several values of $Bo$ (including $Bo$ just below the opening threshold for each $Oh$). At low viscosity ($Oh\ll 0.1$), we find that $t_{c}\sim\tau_{\gamma}$, insensitive to $Oh$. At higher viscosity ($Oh\gg 0.1$), $t_{c}/\tau_{\gamma}\sim Oh$.
These trends can be understood by considering the dominant balance of forces resisting hole closure. For small $Oh$ (inertia-dominated regime), viscous resistance is negligible and inertial forces dominate. Balancing surface tension and inertia yields a characteristic rim velocity $U_{0}\sim\sqrt{\gamma/(\rho R_{0})}$ and thus $t_{c}\sim\tau_{\gamma}\equiv\sqrt{\rho R_{0}^{3}/\gamma}$. In contrast, for large $Oh$ (viscosity-dominated regime), balancing surface tension with viscous stress (of order $\eta U_{0}/R_{0}$) gives $U_{0}\sim\gamma/\eta$ and hence $t_{c}\sim\eta R_{0}/\gamma\sim\tau_{\gamma}Oh$, a visco-capillary timescale. These scaling predictions agree well with the simulation results in fig. 3(a).
Notably, even when $Bo$ is very close to its critical value $Bo_{c}(Oh)$ (beyond which the sheet would open), the measured healing time $t_{c}$ is essentially the same as for much smaller $Bo$. Thus, while radial driving ($Bo$) determines whether the sheet ultimately opens or heals, the healing dynamics themselves is primarily controlled by the liquid’s inertial-viscous balance ($Oh$).
We can predict the boundary between opening and healing regimes by comparing the timescales for rim retraction versus radial sheet advection. For the sheet to open, outward advection (driven by $\omega$, with timescale $t_{\text{adv}}\sim 1/\omega$) must outpace capillary-driven rim closure (timescale $t_{c}$). At low viscosities (small $Oh$, where $t_{c}\sim\tau_{\gamma}$), the criterion $t_{c}\sim t_{\text{adv}}$ leads to $\sqrt{\rho R_{0}^{3}/\gamma}\sim 1/\omega$, yielding $Bo\sim 1$ in dimensionless terms. At high viscosities (large $Oh$, where $t_{c}\sim\eta R_{0}/\gamma$), the condition $t_{c}\sim 1/\omega$ gives $Oh\sim Bo^{-1/2}$. These theoretical thresholds – shown as the gray (vertical) and black (diagonal) lines in fig. 2(c) – closely match the transition observed in our simulations.
What happens in the limiting case of no external driving ($Bo=0$)? – In the absence of radial forcing ($Bo=0$), the liquid sheet inevitably heals for any finite $Oh$ – even extremely viscous sheets (large $Oh$) will eventually reseal, albeit very slowly as viscosity prolongs the rim collision time. This follows from our thresholds: as $Bo\to 0$, achieving opening would require $Oh\sim Bo^{-1/2}\to\infty$, so any finite $Oh$ lies in the healing regime. Our simulations confirm this universal healing at $Bo=0$, consistent with the classic geometric energy criterion of Taylor & Michael [40]. Using a toroidal model for the hole’s rim [41], they predicted that without external driving a circular hole can continue to expand only if its radius exceeds a purely geometric threshold. In particular, when the hole’s outer radius $R$ satisfies $R/h_{0}>\pi/4$ (so that the inner radius $R_{h}=R-h_{0}/2$ is about $0.29h_{0}$), the excess surface energy $\Delta\Xi$ becomes positive, signaling that the hole will grow rather than close. Holes smaller than this critical size have $\Delta\Xi<0$ and will heal. Notably, this threshold is independent of viscosity. Thus, at $Bo=0$ no finite-$Oh$ case can open unless the initial hole is above the critical $R/h_{0}$ – indeed, the cavities in our simulations were below this size, so they all heal.
In more realistic situations, additional physics can delay the initial rupture (for example, due to chemical or thermal heterogeneity), allowing the sheet to drain longer and thin more broadly before the hole forms. We represent such delays in our model by a larger initial cavity distortion characterized by an increased polar angle $\theta$ (fig. 1c). Simulations at $Bo=0$ with varied $\theta$ reveal a clear geometry-controlled threshold: for sufficiently large distortions (large $\theta$), the hole opens even without any driving, whereas for smaller initial distortions the hole heals (fig. 3b). This confirms the second threshold of the double-threshold framework: a sufficiently large initial cavity alone can trigger sheet rupture in the absence of external forcing. Moreover, at very high viscosity (large $Oh$), the opening-healing transition occurs at a nearly constant $\theta\approx 0.09\pi$, consistent with the $R_{h}/h_{0}\approx 0.29$ criterion seen in our simulations at large $Oh$. As $Oh$ decreases (inertia-dominated regime), this critical distortion $\theta_{c}$ rises, since added inertia aids rim closure—requiring a larger initial opening to overcome healing. Remarkably, sheets that do open exhibit non-monotonic cavity growth (fig. 3c). Surface tension first drives the cavity edges apart, causing radial cavity growth, but the induced inertia pulls the rim inward and temporarily reduces the cavity radius. However, the cavity remains large enough to overcome healing by this recoil and eventually opens with the Taylor–Culick mechanism.
Figure 3: (a) After the bubble cavity opens, in some cases, capillarity manages to heal the rims, and the time taken is referred to as the collision time $t_{c}$. Here, $t_{c}/\tau_{\gamma}$ is plotted against OhOh at several BoBo, where $\tau_{\gamma}=\sqrt{\rho R_{0}^{3}/\gamma}$. At small OhOh, the gray line shows the scaling of the inertio-capillary timescale, while the black line shows the visco-capillary time scale at large OhOh; both scalings seem to be consistent with simulation results. (b) The parameter space of $Oh-\theta$ highlighting the healing and opening regimes with different colors. The gray line shows the transition observed at $\theta=0.09\pi$ at large OhOh, while the individual data points are also denoted. (c) The evolution of minimum tip radius $r_{\text{min}}$ at several initial distortions $\theta$ for $Oh=0.1$ without external driving. For smaller distortions, $r_{\text{min}}$ decays to zero, whereas for larger distortions, sheets eventually open irreversibly, consistent with Taylor-Culick retractions. At several instances, the insets depict sheet profiles and highlight $r_{\text{min}}$ with gray arrows.
Effect of off-center bubbles – Thus far we considered sheet-centered bubbles ($\chi/R_{0}=0$) with simultaneous holes at both poles; in practice, an off-center bubble ($\chi/R_{0}\neq 0$) breaks this symmetry. For $\chi/R_{0}>0$, only the thinner pole ruptures initially, leaving a draining liquid bridge at the opposite pole. If the offset is very small ($\chi\to 0$), that remaining bridge is extremely thin and often ruptures almost immediately due to van der Waals forces – essentially the symmetric outcome. For moderate asymmetry, however, the first opening launches a capillary wave that travels toward the far pole and drives fluid into the intact bridge (fig. 4), temporarily replenishing it and suppressing prompt hole nucleation there – unlike the symmetric case (fig. 4a).
Figure 4: The draining liquid sheet of the bubble is placed asymmetrically with $\chi/R_{0}=0.1$. Time evolution has been shown for cases with $Oh=1$, (a) $Bo=10^{-4}$, and (b) $Bo=10^{-2}$. In the former case, the liquid bridge at the south pole replenishes and grows into a flat-shaped sheet due to capillary action, and then the sheet keeps on thinning due to radial drainage. Meanwhile, in the latter case, thinning due to radial drainage dominates before the bridge can be replenished by the damped capillary waves at large $Oh=1$. The left panel depicts velocity magnitude $\|\bm{v}\|/V_{0}$, where $V_{0}=\sqrt{Bo\gamma/\rho R_{0}}$, the black arrows depict the velocity direction. The right panel depicts viscous dissipation $\xi=2Bo\left(\bm{\mathcal{D}:\mathcal{D}}\right)$, where $\bm{\mathcal{D}}=\left(\bm{\nabla u+\nabla u}^{T}\right)/2$ is the symmetric part of velocity gradient tensor. (c) Sheet breakup time (normalized by the no-bubble case rupture time $\tau^{*}\sim\tau_{\gamma}/\sqrt{Bo}$) in the OhBoOh-Bo parameter space for $\chi/R_{0}=0.1$. The gray line depicts the transition line observed for symmetric cases ($\chi/R_{0}=0$), as shown in fig. 2.
[…]
Conclusion & Outlook – We have shown that a micron-thick draining sheet pierced by a trapped bubble undergoes irreversible rupture only if a double-threshold is exceeded: the driving (e.g., airflow or inertia) must exceed a critical Bond number $Bo_{c}(Oh)$, and the initial cavity distortion – captured by a geometric opening angle $\theta$ – must exceed a threshold $\theta_{c}(Oh)$. If either threshold is unmet, capillarity heals the sheet on an inertio- ($Oh\ll 1$) or visco-capillary ($Oh\gg 1$) timescale. This double-threshold mechanism is reminiscent of hole formation in drops impacting on rough surfaces, where a sufficiently high inertia and a large surface roughness are both required for holes to nucleate [17, 13, 46]. Similar double-thresholds govern other nonlinear transitions in fluids: for example, turbulence in shear flow develops only when the Reynolds number is high and the initial perturbation amplitude is large [47], and elastic turbulence in polymer solutions likewise demands a large Weissenberg number together with strong disturbances [48]. Importantly, an isolated flat liquid sheet in air is linearly stable as small distortions decay and rupture only occurs through finite amplitude distortions [49, 50, 51, 52, 53]. This nonlinearity of the healing–opening transition is reflected in our simulations, where small initial cavities decay while sufficiently large ones rupture (fig. 3c). However, unlike the subcritical double-threshold mechanism behavior in transition to turbulence, our system is not subcritical and thus not hysteretic: once parameters ($Bo$, $\theta$, $Oh$) are set, the outcome (healing or opening) is uniquely determined.
A trapped bubble – acting as a hydrophobic defect – thus rationalizes why rupture occurs far above molecular scales and provides predictive control knobs, namely driving strength and defect geometry, relevant to bag-breakup sprays and respiratory films. Future studies could test whether analogous thresholds govern rupture triggered by solid particles, oil droplets, or Marangoni-driven inhomogeneities. Chemical or thermal gradients may couple nonlinearly with radial drainage and shift $Bo_{c}(Oh)$ and $\theta_{c}(Oh)$, extending this double-threshold framework to a broader class of impurity-triggered, geometry-sensitive instabilities.
Acknowledgements.
Acknowledgments: We acknowledge the funding from the MIST consortium with project number P20-35 of the research programme Perspectief, which is (partly) financed by the Dutch Research Council (NWO). We also acknowledge the NWO-Canon grant FIP-II grant. This work was carried out on the national e-infrastructure of SURFsara, a subsidiary of SURF cooperation, the collaborative ICT organization for Dutch education and research. This work was sponsored by NWO - Domain Science for the use of supercomputer facilities. Stéphane Zaleski and Chunheng Zhao were funded by the European Research Council (ERC) through the European Union’s Horizon 2020 Research and Innovation Programme (Grant Agreement No. 883849 TRUFLOW).